mikewwe13:

Use a graphing tool to solve the system. {3x − 8y = 29 {3x + y = −2 Which ordered pair is the best estimate for the solution to the system? A. (0,−3.75) B. (0, −2) C. (0.75, −2.25) D. (0.5, −3.5)

mikewwe13:

@Vocaloid

Vocaloid:

you are given a graphing tool so simply graph both equations and find the intersection re-writing both equations in y = mx + b form may be helpful

mikewwe13:

I just solved it, The answer is either B or D

Vocaloid:

show me the graph you have made

mikewwe13:

well my phone cord is properly working at the moment but my answer shows the appropriate answer so i think it is A

mikewwe13:

not

mikewwe13:

sorry i left that out

Vocaloid:

look at your graph and tell me the x- and y-coordinates of the point

mikewwe13:

the x coordinate is -1 and y coordinate is 2

Vocaloid:

then your graph is incorrect 3x − 8y = 29 3x + y = −2 rewriting this gives us y = (3/8)x - 29/8 y = -3x - 2 graph these two lines and find the intersection please be honest with me if you are having trouble with the graph *** just had to correct a typo, it was - 29/8 not + 29/8

mikewwe13:

1 attachment
mikewwe13:

i have a pic, i finally got it to work

mikewwe13:

the answer is B

Vocaloid:

can you try graphing it again, I made a typo in the original problem just change the + 29/8 to a -29/8

mikewwe13:

OH OK

mikewwe13:

1 attachment
mikewwe13:

there

mikewwe13:

ok now the answer is D

Vocaloid:

good, well done